OpenStudy (mysticfury):

Can somebody please help me with 5 questions? i really need these done. Questions are posted below

OpenStudy (mysticfury):

Which is an equation of the given line in standard form? A. –8x + 9y = 23 B. –8x + 9y = –23 C. –8x + 7y = 25 D. –9x + 8y = –23 https://static.k12.com/calms_media/media/1331000_1331500/1331297/1/c5c8e45fd70b35572ab9141472538d3ecb31705d/HS_AL1G2_S1_07_10_ART2.jpg

OpenStudy (airjordan200):

what do u think it is first

OpenStudy (mysticfury):

i have no idea

OpenStudy (mysticfury):

can you help @airjordan200 ?

OpenStudy (airjordan200):

oh sorry i was in my own world

OpenStudy (mysticfury):

its alright

OpenStudy (mysticfury):

@Atsie do you know how to do this?

OpenStudy (airjordan200):

its C

OpenStudy (mysticfury):

https://static.k12.com/calms_media/media/1312500_1313000/1312644/1/4588fdf88106392accde75ab89d0403ed4be21f7/HS_A1G2_S1_07_10_ART4.jpg Which is the equation of the given line? A. x = –4 B. y = –4 C. y = –4x D. x = 4

OpenStudy (mysticfury):

how about this one?

OpenStudy (mysticfury):

i would say d since there both negative

Atsie (atsie):

I think D would be correct @MysticFury

OpenStudy (mysticfury):

thanks @Atsie

Atsie (atsie):

Sure, no problem :)

OpenStudy (mysticfury):

https://static.k12.com/calms_media/media/308500_309000/308808/1/c9ddd235e76faa5f12cc74a55ca077ec292fb734/MS_ALG_S1_07_10_quiz_Q4_question.gif Which is the equation of the given line in slope-intercept form? A. y = -3x - 1 B. y = 3x - 1 C. y = -1/3x - 1 D. y = 3x + 1

OpenStudy (mysticfury):

D maybe?

Atsie (atsie):

Hmm, I'm not familiar with the how to of slope intercept form so I'm not sure. Did you learn how to do it?

OpenStudy (mysticfury):

@pooja195 taught me how to turn things into slope intercept form but idk how to determine what the equation of the line is to turn it into slope intercept form

OpenStudy (anonymous):

dude don't trust @airjordan he might lie to u

OpenStudy (mysticfury):

@savana24 well at the moment any answer is a good answer, and @airjordan200 has helped me before so i trust him a bit

Atsie (atsie):

Grr, I'm sorry that I can't help you there. I'm completely lost on slope intercept type stuff. I still have yet to learn about it myself. :/

OpenStudy (anonymous):

ok but still i dated him so i wouldn't right now he is in a bad mood

OpenStudy (mysticfury):

its alright @Atsie you helped me on the other one :D

OpenStudy (mysticfury):

that must be why hes not replying @savana24

OpenStudy (airjordan200):

no i was on fb

OpenStudy (mysticfury):

@savana24 you do have any idea what this one could be?

Atsie (atsie):

Maybe you should tag more people. Mathmale is online.

OpenStudy (mysticfury):

oh welcome back @airjordan200

OpenStudy (mysticfury):

ok @mathmale , @pooja195 , @ash_heather1 , can anyone help?

OpenStudy (anonymous):

i am so sorry i am not good at math, and airjordan200 u don't have a FB so stop lying to people

OpenStudy (mysticfury):

maybe he got one?

Atsie (atsie):

Okay, let me just say this right now... @savana24 and @airjordan200 take your personal problems some place else. This is math and last time I check @MysticFury wants help, not two people bickering.

OpenStudy (mysticfury):

imma open this in a new question

Atsie (atsie):

Good idea :) @MysticFury

OpenStudy (anonymous):

no cuz i was just talking to him and looked up his FB account and it said no results of whatever and he is 11 years old so who would let him get a FB

OpenStudy (mysticfury):

lots of kids get FB younger then that

OpenStudy (anonymous):

ok whatever

OpenStudy (anonymous):

caleb i am waiting for u to prove it

OpenStudy (anonymous):

im so sorry i have no clue :/

OpenStudy (anonymous):

@airjordan200 did u get my message in the chat room